OpenStudy (anonymous):

x+5y=39 -x+9y=59 elimination method

OpenStudy (anonymous):

after eliminating 14y=98 y=7 x+5(7)=39 x+35=39 x=4

OpenStudy (anonymous):

all you do is add the two equations down the columns so the x - x would give zero and the 5y +9y gives the 14y

OpenStudy (anonymous):

so how do you get x?

OpenStudy (anonymous):

since now you have the y value you can put it in either of the original equations, then the only variable you have left is x and solve for it

OpenStudy (anonymous):

can you help me with this one and let me know if I do this correctly? 5x+5y=-11 7x-3y=13 x=12 y=8 5(12)+5(8)=100 7(12)-3(8)=60

OpenStudy (anonymous):

what you have to do first is get it down to one equation but getting either the x's in the two equations or the y's in the two equations gone. for this you would need to multiply the equations to get one equal. so for this one. lets eliminate the x's 7(5x+5y= -11) -5(7x-3y=13) now multiply those out and show me what you get

OpenStudy (anonymous):

it is usually easiest to eliminate by multiplying an equation by the other equations coefficient. like with 5x and 7x, multiply the 5x by 7 and 7x by 5, that way you know they will equal each other, but one has to be negative the other equation so that when the two equations are added the variable gets turned to zero